OpenStudy (anonymous):

4sqrt{x} times 5sqrt{7x} 4sqrt{x} times 5sqrt{7x} @Mathematics

OpenStudy (anonymous):

\[\sqrt[4]x \times \sqrt[5]7x\]

OpenStudy (turingtest):

\[\sqrt[4]{x}*\sqrt[5]{7x}=x^{1/4}(7x)^{1/5}=x^{1/4}(7^{1/5})x^{1/5}=7^{1/5}x^{1/4+1/5}=7^{1/5}x^{9/20}\]that's 7^(1/5)x^(9/20) for the answer (a bit difficult to read the 9)

OpenStudy (turingtest):

oh, is the x under the radical or not?

OpenStudy (turingtest):

if is is under the radical the answer is\[7^{1/5}x^{9/20}=\sqrt[5]7*\sqrt[20]{x^9}\]

OpenStudy (anonymous):

yes it is

OpenStudy (turingtest):

then the answer is above

OpenStudy (turingtest):

do you follow?

OpenStudy (anonymous):

yes thhanks!